OpenStudy (anonymous):

Find the derivative, f(x) = 5x + 9 at x = 2

OpenStudy (anonymous):

that would be f'(x) = 5 so f'(2) = 5

OpenStudy (anonymous):

f'(x)=5 f'(2)=5

OpenStudy (anonymous):

Thank you but how do I get this?

OpenStudy (anonymous):

5x+9 find the derivatives separately the derivative of a constant number is always 0 for 5x use the derivative formula for multiplications the derivative of 1st multiplied by 2nd variable plus the 1st variable multiplied by the derivative of the 2nd

OpenStudy (anonymous):

I'm sorry, but could you type the numbers out and go over it with me?

OpenStudy (anonymous):

5x 1st variable is 5, what is the derivative 2nd variable is x, what is the derivative

OpenStudy (anonymous):

I don't know, what is a derivative?

OpenStudy (anonymous):

\[f'(x)=\lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (anonymous):

does it make any sense

OpenStudy (anonymous):

for example derivative of 8x is 8 , 1/4x=1/4, \[\sqrt{2}x=\sqrt{2}\]

OpenStudy (anonymous):

okay yeah, I get it